OpenStudy (anonymous):

What is the distance between the points (–5, –9) and (–5, 13)? A: 4 B:11 C:16 D:22

OpenStudy (anonymous):

distance formula \[\large d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] Where (x1,y1) and (x2,y2) are your first and second coordinates respectively.

OpenStudy (anonymous):

So it's (-9,13) and (-5,-5)?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I don't know I'm confused.

OpenStudy (anonymous):

Just plug in the numbers into the formula.

OpenStudy (anonymous):

You get a single number as an answer.

OpenStudy (anonymous):

so I multiply -9 and 13, then -5 and -5, then subtract the two numbers?

OpenStudy (anonymous):

Your making this much more complicated than it is lol

OpenStudy (anonymous):

\[d = \sqrt{(-5+5)^2+(13+9)^2}\]

OpenStudy (anonymous):

\[d = \sqrt{(0)^2+(22)^2}\]

OpenStudy (anonymous):

\[d = 22 \]

OpenStudy (anonymous):

I see lol I do tend to make things a lot harder on myself. thank you for the help! I wrote down the formula for the next problems

OpenStudy (anonymous):

Np :)